# latexOverwriting file

### v1

@@ 5 Extensions and perspectives for future work
This section is organized as follows. First, in Subsection 5.1, we present an alternative approach based on stopping times for proving our main results Theorems 2.6 and 2.12. Then, in the following subsections, we discuss three natural extensions of our results. In Subsection 5.2, we show how our mean-field limit result for CBS (Theorem 2.12) can be improved when the initial condition $\overline{\rho}_{0}$ is Gaussian. In Subsection 5.3, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, under which the improved stability estimates in Corollary 3.3 are no longer valid. Finally, in Subsection 5.4, we show how Theorems 2.6 and 2.12 can be generalized to smaller values of $p\geqslant 2$.
@@ 5.1 Proving mean-field results using stopping times
In this section, we present an alternative approach for proving Theorem 2.6 and Theorem 2.12. This approach is based on stopping times, and inspired by a method employed in [33] to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method. A similar approach was employed in [36] to obtain a non-quantitative mean-field estimate. Before presenting the proof, we briefly discuss the main differences between this approach and the indicator set approach presented in Section 2.
@@
    In certain contexts, however, such as the one discussed in Subsection 5.2, being able to exclude a time-dependent event is very useful.
  • The stopping time approach presented hereafter does not enjoy this flexibility, but is convenient in settings where moments of both $\mu^{J}$ and $\overline{\mu}^{J}$ need to be controlled in order to apply the stability estimates obtained in Proposition 3.1 for the weighted moments. This is illustrated in Subsections 5.3 and 5.4.
@@
By Doob’s optional stopping theorem [33, Theorem 3.3], see also [20, Equation 2.29, p.285], the stopped process $(M_{t\wedge\theta_{J}})_{t\geqslant 0}$ is a martingale, with a quadratic variation process given by the stopped quadratic variation process of $M_{t}$, i.e. by $(\langle M\rangle_{t\wedge\theta_{J}})_{t\geqslant 0}$. Therefore, by the Burkholder–Davis–Gundy inequality, we have for all $t\in[0,T]$ that
@@
$\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\big{\rvert}^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}\big{\rvert}^{q}\right]+\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}\overline{X}^{(t)}_{j}\big{\rvert}^{q}\right]\right)\leqslant 2^{q}\kappa.$
@@ Conclusion
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{2(q-p)}{q}}\right),$
@@
5.2 Improving the rate of convergence for Gaussian initial conditions.
It is proved in [11] that, if the initial condition to the mean-field equation 1.5 associated with CBS is Gaussian, then at each time $t$ the solution $\overline{\rho}_{t}$ to the equation is also Gaussian. In this setting, the probability measure $\overline{\rho}_{0}$ has moments of all orders. Using a concentration inequality for Gaussian measures (see [43] and [42, Section 2]), it is possible to prove that the probability of the event
$\Omega_{J,t}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\lvert\overline{X}^{(j)}_{t}(\omega)\rvert^{p}\geqslant R\right\}$
decreases in fact as $\exp\bigl{(}-CJ^{\frac{p}{2}}\bigr{)}$, for a constant $C$ depending on $R$ and on the parameters of the Gaussian distribution. Carrying out the proof of Theorem 2.6 with the sharper bound $\exp\bigl{(}-CJ^{\frac{p}{2}}\bigr{)}$ instead of $CJ^{-2}$ for $\mathbf{P}(\Omega_{J,t})$ as in Lemma 2.5, we obtain for all sufficiently large $p$ a mean-field estimate of the form
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
@@ 5.3 Handling the case where $u>\ell>0$ in Assumption H
    where we used that $\mathbf{P}[A+B\geqslant k]\leqslant\mathbf{P}[A\geqslant k/2]+\mathbf{P}[B\geqslant k/2]$ for any two real-valued random variables $A$ and $B$, because $\{A+B\geqslant k\}\subset\{A\geqslant k/2\}\cup\{B\geqslant k/2\}$. A similar trick, of bounding the probability of an event associated with a stochastic system in terms of a similar event associated with a simpler system and the discrepancy between the two systems, is employed in [19, Theorem 2.10]. The probability $\mathbf{P}[\theta_{J}\leqslant T]$ can then be bounded as follows:
@@
5.4 Removing the lower bound on $p$
In Theorems 2.6 and 2.12, we assume that $p$ is bounded from below by $p_{\mathcal{M}}$ and $2p_{\mathcal{M}}$, respectively. It is possible to replace this restriction by the less restrictive inequality $p\geqslant 2$ in the case of Theorem 2.6, and $p>2$ in the case of Theorem 2.12, provided that the initial condition $\overline{\rho}_{0}$ has sufficiently many moments. We motivate this statement hereafter for the specific case of CBO. Suppose for simplicity that $\overline{\rho}_{0}$ has finite moments of all orders. Then by Theorem 2.6, for $f\in\mathcal{A}(s,\ell,\ell)$ there exists an exponent $p_{\star}\geqslant 2$ such that for all $p\geqslant p_{\star}$ and all $\varepsilon>0$, it holds for some constant $C$ independent of $J$ that
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-(2-\varepsilon)}.$ (5.8)
  • First, note that since
    $\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{r}\right]\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{r}{p}}$
    for any $1\leqslant r\leqslant p$, the existence of a mean-field result of the type (5.8) with $p=p_{\star}$ immediately implies a similar result for all $r\in[2,p_{\star}]$, but with a rate that is degraded by a factor $\frac{r}{p_{\star}}$. In the next bullet point, we present a better approach.
  • Fix $r\in[2,p_{\star}]$. From the proof of Proposition 3.1, it is easily observed that there is $\varrho$ sufficiently large to ensure that for all $R>0$, it holds that
    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\times\mathcal{P}_{\varrho,R},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert+\Bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}\leqslant L\,W_{r}(\mu,\nu),$ (5.9)
    for some constant $L=L(R,r)>0$, see Lemma A.6. In order to obtain a mean-field result for the exponent $r$, we can employ the same approach as in the proof in Subsection 5.1 with $p$ replaced by $r$, but with the stopping time
    $\theta_{J}:=\overline{\tau}_{J}\wedge\tau_{J},\qquad\left\{\begin{aligned} \tau_{J}(R)&=\inf\Bigl{\{}t\geqslant 0:\mu_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}},\\ \overline{\tau}_{J}(R)&=\inf\Bigl{\{}t\geqslant 0:\overline{\mu}_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}}.\end{aligned}\right.$
    for some $p\geqslant p_{\star}\vee\varrho$. The proof then goes through as before, except that we use (5.9) instead of Corollary 3.3, and that the probability $\mathbf{P}[\theta_{J}\leqslant T]$ is bounded from above differently. Reasoning as in (5.7), it is simple to show that
    $\mathbf{P}[\theta_{J}\leqslant T]\leqslant 2\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right]+\frac{2^{p}}{R}\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\right].$
    The first term on the right-hand side decreases as $\mathcal{O}(J^{-2})$ by Lemma 2.5, whereas the second term decreases as $\mathcal{O}(J^{-(2-\varepsilon)})$ by (5.8). This enables to prove that
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{r}\right]\leqslant CJ^{-\min\left\{\frac{r}{2},2-\varepsilon\right\}}.$ (5.10)
@@ A.1 Stability estimates for ratios of integrals
Let us first generalize [10, Lemma 3.1] by showing that there exists a constant $\eta=\eta(p,R)>0$ such that
@@ A.2 Bound on weighted moments
for a constant $C$ depending only on $\beta$ and $f$. When $f$ is unbounded, proving an inequality of this type is more involved. The following is a generalization of [10, Lemma 3.3].
@@
is a non-increasing function of $\gamma$. This follows from [11, Lemma A.3], and could also be proved by showing that $F^{\prime}(\gamma)$ is non-positive for all $\gamma\geqslant 0$. A simple calculation indeed gives that

### v2

@@ 5 Extensions and perspectives for future work
This section is organized as follows. First, in Subsection 5.1, we present an alternative approach based on stopping times for proving our main results Theorems 2.6 and 2.12. Then, in Subsection 5.2, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, under which the improved stability estimates in Corollary 3.3 are no longer valid.
@@ 5.1 Proving mean-field results using stopping times
In this section, we present an alternative approach for proving Theorem 2.6 and Theorem 2.12. This approach is based on stopping times, and inspired by a method employed in [MR1949404] to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method. A similar approach was employed in [MR4553241] to obtain a non-quantitative mean-field estimate, and more recently in [vaes2024sharppropagationchaosensemble] to prove sharp propagation of chaos estimates for an ensemble sampler related to the ensemble Kalman filter [EKS-1-Hofmann-Stuart-Li-GI-2019, MR4123680]. Before presenting the proof, we briefly discuss the main differences between this approach and the indicator set approach presented in Section 2.
@@
    In certain contexts, however, being able to exclude a time-dependent event could be very useful.
  • The stopping time approach presented hereafter does not enjoy this flexibility, but is convenient in settings where moments of both $\mu^{J}$ and $\overline{\mu}^{J}$ need to be controlled in order to apply the stability estimates obtained in Proposition 3.1 for the weighted moments. This is illustrated in Subsection 5.2.
@@
By Doob’s optional stopping theorem [MR1949404, Theorem 3.3], see also [MR0838085, Equation 2.29, p.285], the stopped process $(M_{t\wedge\theta_{J}})_{t\geqslant 0}$ is a martingale, with a quadratic variation process given by the stopped quadratic variation process of $M_{t}$, i.e. by $(\langle M\rangle_{t\wedge\theta_{J}})_{t\geqslant 0}$. Therefore, by the Burkholder–Davis–Gundy inequality, we have for all $t\in[0,T]$ that
@@
$\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\big{\rvert}^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}\big{\rvert}^{q}\right]+\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}\overline{X}^{(j)}_{t}\big{\rvert}^{q}\right]\right)\leqslant 2^{q}\kappa.$
@@ Conclusion
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{2(q-p)}{q}}\right),$
@@ 5.2 Handling the case where $u>\ell>0$ in Assumption H
    where we used that $\mathbf{P}[A+B\geqslant k]\leqslant\mathbf{P}[A\geqslant k/2]+\mathbf{P}[B\geqslant k/2]$ for any two real-valued random variables $A$ and $B$, because $\{A+B\geqslant k\}\subset\{A\geqslant k/2\}\cup\{B\geqslant k/2\}$. A similar trick, of bounding the probability of an event associated with a stochastic system in terms of a similar event associated with a simpler system and the discrepancy between the two systems, is employed in [2023arXiv230503637D, Theorem 2.10]. The probability $\mathbf{P}[\theta_{J}\leqslant T]$ can then be bounded as follows:
@@ A.1 Stability estimates for ratios of integrals
Let us first generalize [carrillo2018analytical, Lemma 3.1] by showing that there exists a constant $\eta=\eta(p,R)>0$ such that
@@ A.2 Bound on weighted moments
for a constant $C$ depending only on $\beta$ and $f$. When $f$ is unbounded, proving an inequality of this type is more involved. The following is a generalization of [carrillo2018analytical, Lemma 3.3].
@@
is a non-increasing function of $\gamma$. This follows from [CBS-Carrillo2021, Lemma A.3], and could also be proved by showing that $F^{\prime}(\gamma)$ is non-positive for all $\gamma\geqslant 0$. A simple calculation indeed gives that